How would you make 600 uL of DNA at a concentration of 60 ng/ul from stock DNA that has a concentration of 300 ng/ul and TE buffer?
Answer #1

In this situation, to prepare 600 ul of 60 ng/ul DNA concentration solution from stock of 300 ng/ul, we need to add the volume, x, (say),

Using dilution method,

=> M1V1 = M2V2

=> V2 = M1V1/M2 = 60 * 600/300 = 120 ul

Therefore, to prepare required solution, take 120 ul from the stock and dilute it up to 600 ul.